OpenStudy (anonymous):

how do you fine the dy/dx for the function e^{−x{^3}}?

OpenStudy (anonymous):

What is the rule for the derivative of e^x?

OpenStudy (goformit100):

d [e^{−x{^3}}]/dx

OpenStudy (anonymous):

d/dx e^x =e^x

OpenStudy (campbell_st):

use the rule for \[y = e^{f(x)}\] then the derivative is \[\frac{dy}{dx} = f'(x) \times e^{f(x)}\]